Data Set:

Ruby eyes, curved wings, males - 124

Curved wings, males - 127

wild type, females - 825

wild type, males - 412

Ruby eyes, males - 418

Curved wings, females - 261

Conduct a chi-square test, ensuring to include genetic symbols and genetics hypothesis for the genes. Im not sure where to start or how to determine the P, F1 and F2 generations. Thank you for your help in advance!
Answer #1

Chi square test

X^2=total (observed-expected value)/expected value.

Ruby eyes, curved wings, males observed- 124

Expected-125

Curved wings, males observed- 127 ,expected-120

wild type, females observed- 825,expected-830

wild type, males observed - 412,expected-410

Ruby eyes, males observed- 418,expected-420

Curved wings, females observed - 261,expected-260

In this case work outs are shown as:

(124/125)^2/125+(127-120)^2/120+(412/410)^2/410+(418/420)^2/420+(261-260)^2/260+(825-830)^2/830

=0.131

In 4types male and female are there so degree of freedom=4-1=3

The value is 6.251